OpenStudy (anonymous):

sin^2 theta - cos^2 theta = -1 i know the answer is pi over 4, but how do i get there?

zepdrix (zepdrix):

\[\Large\rm \sin^2\theta-\cos^2\theta=-1\]Multiply both sides by -1,\[\Large\rm \cos^2\theta-\sin^2\theta=1\]Applying Cosine Double Angle Formula,\[\Large\rm \cos2\theta=1\]Leading to,\[\Large\rm 2\theta=0+2k \pi\]Divide by 2 to finish it up. Hmm the answer is not pi/4, I'm not sure where you got that from :[

myininaya (myininaya):

or recall sin^2(theta)+cos^2(theta)=1 then add the two equations/ together

myininaya (myininaya):

oops 25 days old my bad...

zepdrix (zepdrix):

Haha! Did it pop up on the list somehow? :)

myininaya (myininaya):

It was high up there I think. I don't think I did too much scrolling. :p